# Leverage Knowledge Graph and Large Language Model for Law Article Recommendation: A Case Study of Chinese Criminal Law

### v2

@@ Abstract
Court efficiency is vital for social stability. However, in most countries around the world, the grassroots courts face case backlogs, with decisions relying heavily on judicial personnel’s cognitive labor, lacking intelligent tools to improve efficiency. To address this issue, we propose an efficient law article recommendation approach utilizing a Knowledge Graph (KG) and a Large Language Model (LLM). Firstly, we propose a Case-Enhanced Law Article Knowledge Graph (CLAKG) as a database to store current law statutes, historical case information, and correspondence between law articles and historical cases. Additionally, we introduce an automated CLAKG construction method based on LLM. On this basis, we propose a closed-loop law article recommendation method. Finally, through a series of experiments using judgment documents from the website ”China Judgements Online”, we have improved the accuracy of law article recommendation in cases from 0.549 to 0.694, demonstrating that our proposed method significantly outperforms baseline approaches.
@@ 1 Introduction
In a modern rule-of-law society, the efficiency of court judgments is crucial for ensuring economic growth and social stability[1, 2, 3]. Meanwhile, [4] revealed that the current backlog of cases in courts is enormous, which means the workload of judicial personnel is heavy. [5] showed that there is a great need for intelligent tools to assist judicial personnel in improving the efficiency of case adjudication, since intelligent tools boost judicial efficiency by categorizing similar cases, speeding up research, and staying updated with new laws, ensuring faster and more consistent decisions.
@@
Law article recommendation entails predicting the relevant law articles applicable to a case based on its factual description. With the advancement of computer science, many data science techniques have been applied to this task, greatly improving efficiency. Classic law article recommendation techniques, such as LSTM[6], Text-CNN[7], GRU[8], RCNN[9], HAN[10], and DPCNN[11], are based on traditional machine learning methods. These methods focus solely on the correspondence between case facts and law article ids, while overlooking the semantic information in law articles. Moreover, these methods are highly susceptible to the effects of insufficient data and imbalanced data labels. Therefore, there is a need to solve above problems.
@@
Some researchers utilized Large Language Models (LLMs) to complete the task of law article recommendation[12, 13, 14]. By using a large LLM, users can not only retrieve the law article in numeric and textual form, but also investigate the rationale behind its applicability to the case by asking follow-up questions. Moreover, users have the capability to probe into the rationale for the non-selection of other analogous provisions. However, [15] found that directly using LLMs is prone to hallucinations, generating incorrect law article information. Therefore, it is necessary to introduce a fully accurate legal information and case knowledge base for auxiliary analysis.
@@
To mitigate the issue of hallucination, researchers introduced Retrieval-Augmented Generation (RAG) technology to enhance the generation capability of LLMs, by retrieving relevant information from external knowledge bases[16, 17]. Currently, RAG technology has been applied in fields such as code generation[18], autonomous driving[19], and enterprise management[20]. However, research on RAG methods in the domain of law article recommendation remains relatively underexplored. RAG is divided into two parts: retrieval and generation. In academic discussions, greater attention is often given to the retrieval component of RAG, whereas the generation process is predominantly handled by LLMs. Classic retrieval methods include TFIDF-based retrieval[21], BM25-based retrieval[22], and retrieval methods based on frozen BERT[23]. However, the above methods mainly focus on the word level and rarely use macro semantics for matching.
@@
To deal with that problem, [24] proposed that the Knowledge Graph (KG), an unstructured database that uses a directed graph structure to store data, is highly suitable to serve as a database for RAG technology. [25] have pointed out that the integration of KGs with LLMs shows great potential for future research. [26] have utilized LLMs to automatically construct knowledge graphs, which are then used to enhance the generation capabilities of LLMs. This method performs well in terms of comprehensiveness and diversity of answers. However, the KGs generated using this method often have complex structures and lack a fixed schema, making them difficult for users to modify, such as adding new cases or removing outdated ones.
To address the above issues, we introduce knowledge graph technology and large language model technology to improve the output accuracy in law article recommendation tasks. Initially, we pre-design the schema of a Case-Enhanced Law Article Knowledge Graph (CLAKG) capable of simultaneously storing current law articles and historical cases information as a database. On this basis, we also introduce an automated CLAKG construction method based on LLM. Furthermore, we design a closed-loop human-machine collaboration method to complete the law article recommendation task. To demonstrate the effectiveness of our proposed method, we constructed a dataset of Chinese criminal law judgments. We compared the method introduced in this paper with approaches based on BERT[23], GRU[8], DPCNN[11], and several RAG approaches such as TFIDF-RAG[21], Graph-RAG[27] and Light-RAG[28]. Finally, we conducted an ablation study of our method.
@@
The remainder of this paper is organized as follows. Section 2 reviews related works about law article recommendation, RAG technology and knowledge graphs in legal domain. Section 3 proposes a closed-loop law article recommendation framework. Section 4 utilizes the method from the third section to construct the CLAKG and compares the accuracy of this method with other baseline approaches. Lastly, the conclusions and future work are highlighted in Section 5.
@@ 2.1 Law Article Recommendation
Early approaches to law article recommendation treated the task as a text classification or retrieval problem. Classical methods relied on keyword matching or manual feature engineering, but these struggled with the complex language of statutes. With the rise of deep learning, researchers applied neural models like LSTMs[6] and CNNs[7] to match case fact descriptions with relevant law articles. These models improved over bag-of-words baselines by capturing sequential and local context features. Transformer-based language models further advanced the field of BERT[29] and its legal-domain adaptations (e.g., LegalBERT[30]) enabled richer semantic representations of legal text. For example, Chalkidis et al.[31] showed that a BERT-based classifier can predict applicable European human rights articles more accurately than traditional models, while Aletras et al.[32] demonstrated early on the feasibility of data-driven prediction of court decisions and their pertinent articles. Despite these gains, most prior models formulate law article recommendation as mapping facts to article IDs, without directly leveraging the content of the law articles themselves. This focus solely on fact-to-ID correspondence makes them heavily data-dependent and sensitive to label imbalance in training sets[33]. In practice, it also limits interpretability – the models do not explain why a recommended article is relevant, since the law articles are not explicitly used in the reasoning process. These limitations have motivated explorations into techniques that can incorporate external legal knowledge and provide better justification for recommendations.
@@ 2.2 LLMs and Retrieval-Augmented Generation Techniques
The advent of large language models (LLMs), such as GPT-style transformers, has opened new avenues for law article recommendation. Rather than training a specific classifier, one can prompt an LLM with a case’s facts and ask for relevant law articles. Recent works have started to investigate this approach[12, 13, 14], noting that LLMs can, in principle, leverage their vast pretrained knowledge to identify relevant law articles and even generate explanations. In particular, Shui et al.[14] evaluate various state-of-the-art LLMs on legal judgment prediction subtasks (including relevant law article prediction), finding that while LLMs exhibit promising zero-shot capabilities, they still lag behind task-specific models on accuracy. A major concern with directly using LLMs for legal guidance is hallucination: the model may produce plausible-sounding but incorrect or nonexistent legal references. Dahl et al.[15] provide a sobering analysis of this issue, showing that even advanced LLMs frequently fabricate legal citations or statutes, which is unacceptable in high-stakes domains like law. To mitigate such hallucinations, researchers are turning to Retrieval-Augmented Generation (RAG) techniques. RAG combines LLMs with information retrieval: before or during text generation, the system retrieves relevant documents (e.g. law articles or case precedents) and conditions the LLM on that retrieved evidence. By grounding responses in actual law articles, RAG can significantly improve factual accuracy and trustworthiness of the model’s output. This approach was originally popularized in open-domain QA systems[16] and has since been applied to legal NLP as well. For instance, a RAG-based legal assistant might fetch the text of candidate law articles given a case description, and the LLM then quotes or summarizes those law articles when recommending applicable law, thereby rooting its answer in verifiable sources. Key components of RAG include a retriever and a generator. In the legal domain, classic IR methods like TF–IDF[21] or BM25[22] have been used to retrieve statute candidates by keyword overlap, while more recent approaches use neural retrievers (e.g. dense embeddings). Borgeaud et al.[23] showed that even a frozen BERT-based retriever can substantially improve generation quality by supplying relevant text from a large corpus. However, naive retrieval based on surface text may miss conceptual matches. Legal phrasing varies, and a model might need to retrieve statutes using synonymous terms or implied connections, something neural retrieval aims to handle better than lexical methods. Overall, RAG represent a promising direction for law article recommendation, as they combine the reasoning ability of generative models with the grounding of information retrieval, thereby reducing hallucinations and improving the transparency of the recommendation.
@@ 2.3 Knowledge Graphs in the Legal Domain
Beyond unstructured text retrieval, Knowledge Graphs (KGs) have emerged as valuable resources for legal information retrieval and recommendation. A legal knowledge graph represents entities such as law articles, cases, legal concepts, and their relationships (e.g., a case cites a statute, or a chapter includes certain articles) in a structured form. This graph structure can complement text-based methods by capturing the domain’s ontology and the interconnections between law articles. Prior works have constructed legal knowledge graphs to assist various tasks. For example, law-specific knowledge bases have been used to integrate charge definitions and article relationships into prediction models, yielding performance gains and better interpretability in tasks like judgment prediction[34, 35]. More recently, Chaudhri et al.[24] argued that KGs are well-suited as the underlying database for RAG systems: instead of retrieving free text, one can query a knowledge graph to retrieve facts or linked nodes, ensuring that the generation is grounded in a consistent knowledge structure. In the legal domain, a KG could be queried to find statutes related to certain legal concepts or past cases with similar fact patterns. One advantage of KGs is that they store high-level associations beyond lexical similarity – for instance, a graph can link a section of law to broader categories or to precedent cases, enabling the recommendation system to recall relevant law articles even if a case’s description uses different wording. There is growing interest in integrating KGs with LLMs[25]. Pan et al. [25] outline a roadmap for unifying the strengths of symbolic knowledge graphs and distributed language models, noting that KGs can provide factual grounding for LLMs while LLMs can help populate and expand KGs from raw texts. In the context of law, researchers have begun designing hybrid systems that use KGs alongside LLMs. Chen et al.[33], for instance, construct a case-enhanced law article KG that links statutes with relevant past cases, and then use an LLM-based pipeline to recommend articles by traversing the graph for pertinent nodes (law articles) and using the LLM to generate the final recommendations and explanations. Such approaches have demonstrated improved accuracy over text-only baselines, highlighting that the structured knowledge in KGs can effectively guide the retrieval and generation process. Meanwhile, Edge et al.[26] explore automatic construction of knowledge graphs using LLMs (generating graph triples from unstructured legal text) and using those graphs to enhance answer generation. This suggests a future where legal KGs and LLMs continuously enrich each other: the KG offers a source of truth to reduce LLM errors, and the LLM helps keep the KG up-to-date with new insights. The integration of knowledge graphs in legal AI is still an emerging area, and challenges remain, such as ensuring the completeness and currency of the graph. Nonetheless, the combination of KGs with retrieval-augmented LLMs holds great promise for building more reliable and interpretable law article recommendation systems in the coming years.
@@ 3 Method
Figure 1: The pipeline of CLAKG. The Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG).
The pipeline of CLAKG-based law article recommendation is shown in Figure 1. We first establish a knowledge graph based on the designed schema. Initially, nodes and relationships are extracted from law articles and judgments in cases to construct the Law Article Knowledge Graph (LAKG) and the Adjudicated Cases Knowledge Graph (ACKG). Subsequently, by utilizing a Large Language Model (LLM), these two graphs are integrated to form CLAKG. This part is introduced in Section 3.1. Building upon this, we have developed a human-machine Collaborative law article recommendation framework. Users describe new case information, and the LLM would provide potentially applicable law articles based on keyword matching and graph embedding techniques. Moreover, the system can provide similar historical case information from the CLAKG, followed by providing most relevant law articles that match the new case using LLM. This part is introduced in Section 3.2. Following the above process, the user reviews the law articles provided by the system and corrects any inaccuracies. The new case information is used to update the CLAKG.

### v3

@@ Abstract
Judicial efficiency is critical to social stability. However, in many countries worldwide, grassroots courts face substantial case backlogs, and judicial decisions remain heavily dependent on judges’ cognitive efforts, with insufficient intelligent tools to enhance efficiency. To address this issue, we propose a highly efficient law article recommendation approach combining a Knowledge Graph (KG) and a Large Language Model (LLM). First, we construct a Case-Enhanced Law Article Knowledge Graph (CLAKG) to store current law articles, historical case information, and their interconnections, alongside an LLM-based automated construction method. Building on this, we propose a closed-loop law article recommendation framework integrating graph embedding-based retrieval and KG-grounded LLM reasoning. Experiments on judgment documents from China Judgments Online demonstrate that our method boosts law article recommendation accuracy from 0.549 to 0.694, outperforming strong baselines significantly. To support reproducibility and future research, all source code and processed datasets are publicly available on GitHub (see Data Availability Statement).
@@ 1 Introduction
In a modern rule-of-law society, judicial efficiency is critical to safeguarding economic growth and social stability[12, 25, 17]. Meanwhile, [21] noted that courts currently face a massive case backlog, resulting in a heavy workload for judicial personnel. [43] emphasized an urgent need for intelligent tools to assist judicial personnel in enhancing case adjudication efficiency, as they boost judicial efficiency by categorizing similar cases, accelerating legal research, and keeping abreast of new laws—thus ensuring faster and more consistent decisions.
@@
Law article recommendation involves predicting the relevant law articles applicable to a case based on its factual description. With the development of computer science, numerous data science techniques have been applied to this task, significantly enhancing efficiency. Classic law article recommendation techniques—including LSTM[22], Text-CNN[26], GRU[8], RCNN[28], HAN[46], and DPCNN[24]—are based on traditional machine learning methods. These methods focus exclusively on the correspondence between case facts and law article IDs, while neglecting the semantic information within law articles. Furthermore, these methods are highly vulnerable to insufficient data and imbalanced data labels. Thus, there is an urgent need to address these issues.
@@
Several researchers have employed Large Language Models (LLMs) for law article recommendation[45, 30, 37]. With large LLMs, users can not only retrieve law articles in numerical and textual forms but also explore the rationale for their applicability to cases through follow-up questions. Furthermore, users can probe the rationale for not selecting other analogous law articles. However, [10] found that direct use of LLMs is prone to hallucinations, producing inaccurate law article information. Thus, it is essential to incorporate a fully accurate legal information and case knowledge base for auxiliary analysis.
Beyond improving the underlying knowledge base, a parallel line of work shows that prompt engineering can itself be a powerful lever for steering LLM behaviour and mitigating hallucinations. Recent surveys on hallucinations in LLMs highlight that a substantial portion of erroneous outputs is prompt-sensitive and can be reduced by disciplined prompting strategies, such as explicit uncertainty instructions, decomposition prompts, and verification-style prompts [23, 13]. Within this line of work, chain-of-thought prompting and its self-consistency variant encourage models to externalize intermediate reasoning and aggregate over multiple reasoning paths, which generally leads to more reliable answers on complex reasoning and knowledge-intensive tasks [42, 41]. Building on the same intuition that exposing and checking intermediate reasoning improves faithfulness, verification-oriented prompting schemes such as Chain-of-Verification [15] ask the model to first draft an answer, then generate targeted follow-up questions, answer them independently, and finally revise the initial answer based on the verified evidence, thereby significantly decreasing hallucination rates in open-domain question answering. Complementary techniques like SelfCheckGPT [31] further exploit prompt-based self-correction by sampling alternative completions and detecting contradictions as a black-box mechanism for filtering hallucinated content. Taken together, these studies suggest that prompt engineering provides an essential, model-agnostic axis for hallucination mitigation that naturally complements both architectural advances and RAG-style grounding. Motivated by this perspective, we design CLAKG-grounded prompts with explicit role specifications and strict output constraints, so that the LLM is systematically nudged to rely on verifiable legal evidence rather than fabricating non-existent law articles.
@@
To mitigate hallucination, researchers have introduced Retrieval-Augmented Generation (RAG) technology to enhance LLMs’ generation capabilities by retrieving relevant information from external knowledge bases[29, 4]. Currently, RAG technology has been applied in domains such as code generation[18], autonomous driving[11], and enterprise management[32]. However, research on RAG-based law article recommendation remains relatively underexplored. RAG consists of two components: retrieval and generation. In academic discourse, greater attention is often paid to RAG’s retrieval component, while the generation process is primarily handled by LLMs. Classic retrieval methods include TF-IDF-based retrieval[1], BM25-based retrieval[35], and frozen BERT-based retrieval[3]. However, these methods primarily focus on the lexical level and rarely leverage macro-semantic information for matching.
@@
To address this issue, [7] proposed that Knowledge Graphs (KGs)—structured databases that use directed graph structures to store data—are highly suitable as knowledge bases for RAG technology. [34] noted that integrating KGs with LLMs holds significant potential for future research. [16] employed LLMs to automatically construct KGs, which are then used to enhance LLMs’ generation capabilities. This approach achieves strong performance in terms of answer comprehensiveness and diversity. However, KGs generated via this approach often have complex structures and lack a fixed schema, making them difficult to modify—for example, adding new cases or removing outdated ones.
To address the above issues, we propose to combine knowledge graph technology with large language models to improve the accuracy and interpretability of law article recommendation. We first pre-design the schema of a Case-Enhanced Law Article Knowledge Graph (CLAKG) that can simultaneously store current law articles and historical case information in a task-oriented manner. On this basis, we develop an automated CLAKG construction pipeline driven by LLMs. Furthermore, we build a closed-loop human–machine collaboration mechanism in which CLAKG supports retrieval-augmented prompting for LLMs, while expert feedback is continuously injected back into CLAKG to refine its structure.
To assess the effectiveness of the proposed framework, we focus on Chinese criminal law as a representative civil-law domain and construct a dataset of criminal judgments from China Judgments Online. We compare our method with strong text-classification baselines based on BERT[3], GRU[8], and DPCNN[24], as well as several RAG approaches such as TFIDF-RAG[1], Graph-RAG[16], and Light-RAG[20]. We further conduct ablation studies to investigate the contribution of different components of our framework.
Building on the above design and empirical study, this work makes the following main contributions:
  1. 1.
    We propose a Case-Enhanced Law Article Knowledge Graph (CLAKG) that unifies law articles and adjudicated cases under a task-oriented schema, together with an automated CLAKG construction pipeline based on LLMs.
  2. 2.
    We design an LLM–CLAKG-based law article recommendation framework that integrates CLAKG with LLMs in a retrieval-augmented generation (RAG) architecture. By grounding the LLM in CLAKG during recommendation, the framework mitigates hallucinations.
  3. 3.
    We build a Chinese criminal law dataset and conduct extensive experiments showing that our method improves law article recommendation accuracy from 0.549 (LLM only) to 0.694 (LLM + CLAKG with case information), consistently outperforming strong text-classification baselines (BERT, GRU, DPCNN) and RAG baselines (TFIDF-RAG, Graph-RAG, Light-RAG).
@@
The remainder of this paper is organized as follows. Section 2 reviews related work on law article recommendation, RAG technology, and knowledge graphs in the legal domain. Section 3 presents the proposed closed-loop law article recommendation framework. Section 4 describes the construction of CLAKG and compares the accuracy of the proposed method with other baseline approaches. Lastly, Section 5 concludes the paper and outlines directions for future work.
@@ 2.1 Law Article Recommendation
Early approaches to law article recommendation treated the task as a text classification or retrieval problem. Classical methods relied on keyword matching or manual feature engineering, but these struggled with the complex language of statutes. With the rise of deep learning, researchers applied neural models like LSTMs[22] and CNNs[26] to match case fact descriptions with relevant law articles. These models improved over bag-of-words baselines by capturing sequential and local context features. Transformer-based language models further advanced the field of BERT[14] and its legal-domain adaptations (e.g., LegalBERT[6]) enabled richer semantic representations of legal text. For example, Chalkidis et al.[5] showed that a BERT-based classifier can predict applicable European human rights articles more accurately than traditional models, while Aletras et al.[2] demonstrated early on the feasibility of data-driven prediction of court decisions and their pertinent articles. Despite these gains, most prior models formulate law article recommendation as mapping facts to article IDs, without directly leveraging the content of the law articles themselves. This focus solely on fact-to-ID correspondence makes them heavily data-dependent and sensitive to label imbalance in training sets. In practice, it also limits interpretability – the models do not explain why a recommended article is relevant, since the law articles are not explicitly used in the reasoning process. These limitations have motivated explorations into techniques that can incorporate external legal knowledge and provide better justification for recommendations.
@@ 2.2 LLMs and Retrieval-Augmented Generation Techniques
The advent of LLMs, such as GPT-style transformers, has opened new avenues for law article recommendation. Rather than training a specific classifier, one can prompt an LLM with a case’s facts and ask for relevant law articles. Recent works have started to investigate this approach[45, 30, 37], noting that LLMs can, in principle, leverage their vast pretrained knowledge to identify relevant law articles and even generate explanations. In particular, Shui et al.[37] evaluate various state-of-the-art LLMs on legal judgment prediction subtasks (including relevant law article prediction), finding that while LLMs exhibit promising zero-shot capabilities, they still lag behind task-specific models on accuracy. A major concern with directly using LLMs for legal guidance is hallucination: the model may produce plausible-sounding but incorrect or nonexistent legal references. Dahl et al.[10] provide a sobering analysis of this issue, showing that even advanced LLMs frequently fabricate legal citations or statutes, which is unacceptable in high-stakes domains like law. To mitigate such hallucinations, researchers are turning to Retrieval-Augmented Generation (RAG) techniques. RAG combines LLMs with information retrieval: before or during text generation, the system retrieves relevant documents (e.g. law articles or case precedents) and conditions the LLM on that retrieved evidence. By grounding responses in actual law articles, RAG can significantly improve factual accuracy and trustworthiness of the model’s output. This approach was originally popularized in open-domain QA systems[29] and has since been applied to legal NLP as well. For instance, a RAG-based legal assistant might fetch the text of candidate law articles given a case description, and the LLM then quotes or summarizes those law articles when recommending applicable law, thereby rooting its answer in verifiable sources. Key components of RAG include a retriever and a generator. In the legal domain, classic IR methods like TF–IDF[1] or BM25[35] have been used to retrieve statute candidates by keyword overlap, while more recent approaches use neural retrievers (e.g. dense embeddings). Borgeaud et al.[3] showed that even a frozen BERT-based retriever can substantially improve generation quality by supplying relevant text from a large corpus. However, naive retrieval based on surface text may miss conceptual matches. Legal phrasing varies, and a model might need to retrieve statutes using synonymous terms or implied connections, something neural retrieval aims to handle better than lexical methods. Overall, RAG methods represent a promising direction for law article recommendation, as they combine the reasoning ability of generative models with the grounding of information retrieval, thereby reducing hallucinations and improving the transparency of the recommendation.
@@ 2.3 Knowledge Graphs in the Legal Domain
Beyond unstructured text retrieval, Knowledge Graphs (KGs) have emerged as valuable resources for legal information retrieval and recommendation. A legal knowledge graph represents entities such as law articles, cases, legal concepts, and their relationships (e.g., a case cites a statute, or a chapter includes certain articles) in a structured form. This graph structure can complement text-based methods by capturing the domain’s ontology and the interconnections between law articles. Prior works have constructed legal knowledge graphs to assist various tasks. For example, Chaudhri et al.[7] argued that KGs are well-suited as the underlying database for RAG systems: instead of retrieving free text, one can query a knowledge graph to retrieve facts or linked nodes, ensuring that the generation is grounded in a consistent knowledge structure. In the legal domain, a KG could be queried to find statutes related to certain legal concepts or past cases with similar fact patterns. One advantage of KGs is that they store high-level associations beyond lexical similarity – for instance, a graph can link a section of law to broader categories or to precedent cases, enabling the recommendation system to recall relevant law articles even if a case’s description uses different wording. There is growing interest in integrating KGs with LLMs[34]. Pan et al. [34] outline a roadmap for unifying the strengths of symbolic knowledge graphs and distributed language models, noting that KGs can provide factual grounding for LLMs while LLMs can help populate and expand KGs from raw texts. In the context of law, researchers have begun designing hybrid systems that use KGs alongside LLMs. Edge et al.[16] explore automatic construction of knowledge graphs using LLMs (generating graph triples from unstructured legal text) and using those graphs to enhance answer generation. This suggests a future where legal KGs and LLMs continuously enrich each other: the KG offers a source of truth to reduce LLM errors, and the LLM helps keep the KG up-to-date with new insights. The integration of knowledge graphs in legal AI is still an emerging area, and challenges remain, such as ensuring the completeness and currency of the graph. Nonetheless, the combination of KGs with retrieval-augmented LLMs holds great promise for building more reliable and interpretable law article recommendation systems in the coming years.
@@ 3 Method
Figure 1: Overall closed-loop pipeline of the proposed CLAKG-based law article recommendation system. The Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Cases Knowledge Graph (ACKG).
The overall closed-loop pipeline of the proposed CLAKG-based law article recommendation system is shown in Figure 1. It consists of three main stages: (i) constructing the Case-Enhanced Law Article Knowledge Graph (CLAKG) (Section 3.1); (ii) performing LLM–CLAKG based law article recommendation (Section 3.2), as shown in Figure 4; and (iii) human–machine collaboration, where legal experts review and correct the recommended articles and the CLAKG is updated accordingly (Section 3.3).
We first establish a knowledge graph based on the designed schema. Initially, nodes and relationships are extracted from law articles and judgments in cases to construct the Law Article Knowledge Graph (LAKG) and the Adjudicated Cases Knowledge Graph (ACKG). Subsequently, by utilizing a Large Language Model (LLM), these two graphs are integrated to form CLAKG. This part is introduced in Section 3.1. Building upon this, we have developed a human–machine collaborative law article recommendation framework. Users describe new case information, and the LLM provides potentially applicable law articles based on keyword matching and graph embedding techniques. Moreover, the system can provide similar historical case information from the CLAKG, followed by providing the most relevant law articles that match the new case using LLM (Section 3.2). Following the above process, the legal expert reviews the law articles provided by the system and corrects any inaccuracies. The new case information is used to update the CLAKG, forming a closed-loop human–machine collaboration process whose practical effect is illustrated in Section 3.3.